OpenStudy (anonymous):

How do I do this problem? It's asking me this.. Solve: -1<2-3x<1 The answer can be written in form: A

OpenStudy (anonymous):

a = -1 b = -1/3

OpenStudy (anonymous):

-1<-2<-3x<1 A = -1 b=-1/3

OpenStudy (anonymous):

Can you guys explain to me how you got that?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

it said that x must be greater than A and it also must be less than B so you need to transfer the 2 from 2-3x to any side and devide both side by -3 to eliminate it from x,

OpenStudy (anonymous):

the only thing you need to remember when you will transfer the 2 from 2-3x is that you must stisfy the given A<x<B

OpenStudy (anonymous):

so if we transfer 2 to the left side and devided both sides by -3 it will be: 1< x <-1/3 but 1 is greater than -1/3 how can be x is greater than 1 but less than -1/3 so the only other option is to transfer 2 to the right to get the correct answer.

OpenStudy (anonymous):

I put that answer in and it said it was wrong :( AHHHHH I don't get these !

OpenStudy (anonymous):

the right answer is a = -1 b = -1/3

OpenStudy (ybarrap):

1/3 < x < 1

OpenStudy (anonymous):

I did that so the answer is positive or negative?

OpenStudy (anonymous):

Ahhhh ! Since I got it from it gave me a new problem now!

OpenStudy (anonymous):

Now it gave me 1<-2-5x<4

OpenStudy (anonymous):

the answers were negative.

OpenStudy (ybarrap):

-1<2-3x<1 -1 - 2 < -3x < 1 - 2 -3 < -3x < -1 -3/-1 > -3x/-1 > -1/-1 3 > 3x > 1 3/3 > 3x/3 > 1/3 1 > x > 1/3 or 1/3 < x < 1

OpenStudy (anonymous):

Can you help me with this problem because it gave me a whole new problem now :( ! 1<-2-5x<4

OpenStudy (ybarrap):

For 1<-2-5x<4, Ans. -6/5 < x < -3/5

OpenStudy (anonymous):

-3/5 < x < -6/5

OpenStudy (ybarrap):

Solved similarly: 1<-2-5x<4 1+2<-5x<4+2 3<-5x<6 -3 > 5x > -6 -3/5 > x > -6/5 or -6/5 < x < -3/5

OpenStudy (anonymous):

So which ones A and B?

OpenStudy (anonymous):

Cus I have to put it in like that

OpenStudy (ybarrap):

smallest is typically indicated by the first variable as in (A,B); so A = -6/5 and B = -3/5

OpenStudy (nilankshi):

good

OpenStudy (anonymous):

AHHHH I FREAKING LOVE YOU THANKYOUUUUU :)

OpenStudy (ybarrap):

np

OpenStudy (anonymous):

hello, i can tell u it should be -1<-3x+2<1 -1-2<-3x<1-2 -3<-3x<-1 you derive it then to -3 and the position of < will come back to > because you have derived it to a negative one. so: 1>x>1/3\[\rightarrow\] 1/3<x<1 a=1/3 b=1